OpenStudy (anonymous):

What are the vertical and horizontal asymptotes of y=1+(1/x)+(1/x^2)?

OpenStudy (asnaseer):

do you know what happens at vertical asymptotes?

OpenStudy (asnaseer):

what do you think will be the value of y at a vertical asymptote?

OpenStudy (anonymous):

vertical asymptote is x=-1

OpenStudy (anonymous):

I dont know. thats why im asking. I tried doing the limit as x approaches infinity and i got a horizontal asymptote of 1

OpenStudy (asnaseer):

your horizontal asymptote is correct :)

OpenStudy (asnaseer):

now think about what happens at a vertical asymptote - try drawing a picture and see if you can work it out

OpenStudy (anonymous):

How do you get a vertical asymptote of -1? and the whole point of finding the asymptotes is so that i can draw the graph

OpenStudy (asnaseer):

that is wrong - please ignore it

OpenStudy (asnaseer):

but do you understand what a vertical asymptote represents?

OpenStudy (anonymous):

am right

OpenStudy (asnaseer):

horizontal asymptote is limit of y as x approaches infinity so vertical asymptote is ...

OpenStudy (anonymous):

try to do thn u ll knw

OpenStudy (asnaseer):

another way to look at it is as follows: x tends to plus or minus infinity as the curve approaches a horizontal asymptote

OpenStudy (asnaseer):

now think about what a vertical asymptote could mean

OpenStudy (asnaseer):

you can only find it if you know what it represents - i.e. you need to understand what a vertical asymptote is before you can calculate it for this curve.

OpenStudy (anonymous):

horizontal is 1

OpenStudy (asnaseer):

hint: think about what happens as x becomes smaller and smaller

OpenStudy (asnaseer):

what happens to y?

OpenStudy (anonymous):

0=1+(1/x) thn 1/x=-1 then multiply by x ll be -x=1 then divd by -1 both ll be x=-1

OpenStudy (asnaseer):

@semaameya I can only help you if you respond to the questions.

OpenStudy (anonymous):

so when x becomes smaller, does y approach 1?

OpenStudy (asnaseer):

look at the terms in your equation - they involve terms like:\[\frac{1}{x}\]what happens to the value of 1/x as x becomes smaller and smaller - approaching zero

OpenStudy (anonymous):

Yeah. so when x is approaching 0, 1/x will be 1/0 which is undefined. same thing goes to the 1/x^2. so all we have left will be the first term, which is 1

OpenStudy (asnaseer):

no - think about this - what is: 1/0.0001 = ?

OpenStudy (anonymous):

a really big number.

OpenStudy (asnaseer):

so now what happens with this: 1/0.00000000001 = ? is this even larger?

OpenStudy (anonymous):

Yeah. so y approaches positive infinity?

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

so as x tends to zero, y tends to positive infinity

OpenStudy (asnaseer):

so the vertical asymptote is at x=?

OpenStudy (anonymous):

if y approaches positive infinity, does this mean there is no vertical asymptote?

OpenStudy (asnaseer):

no - it means that there is a vertical asymptote there

OpenStudy (anonymous):

at x=0?

OpenStudy (asnaseer):

just like when x tended to infinity, y tended to 1 meant there is a horizontal asymptote at y=1

OpenStudy (asnaseer):

correct - well done! :)

OpenStudy (anonymous):

Thank you so much!.

OpenStudy (asnaseer):

yw :)